HUBBLE SPACE TELESCOPE DAILY REPORT #2453 PERIOD COVERED: 0000Z (UTC) 09/08/99 - 0000Z (UTC) 09/09/99 Daily Status Report as of 252/0000Z 1.0 OBSERVATIONS SCHEDULED AND ACCOMPLISHED: 1.1 Completed STIS/CCD 8409 (STIS CCD Bias Monitor C8) The Space Telescope Imaging Spectrograph (CCD) was used to make a CCD bias calibration observation. This is to monitor the bias in the 1x1, 1x2, 2x1, and 2x2 bin settings at gain=1, and 1x1 at gain = 4, to build up high-S/N superbiases and to track the evolution of hot columns. The observations were executed as scheduled, and no anomalies were reported. 1.2 Completed STIS/CCD 8408 (STIS CCD Dark Monitor C8) The Space Telescope Imaging Spectrograph (CCD) was used to make a CCD bias calibration observation. This is to monitor the darks for the STIS CCD. The observations were executed as scheduled, and no anomalies were reported. 1.3 Completed WF/PC-2 8443 (WF/PC-2 Cycle-8 Supplemental Darks pt1/3) The WF/PC-2 was used to obtain three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were completed as planned, and no problems were reported. 1.4 Completed Two Sets of STIS/CCD/MA2 7752 (The Distribution and Evolution of Lyman-Alpha Forest Cloud Sizes) The Space Telescope Imaging Spectrograph (MA2) was to be used to make ultraviolet observations of the quasar Q0107-025B in order to study the distribution and evolution of the Lyman-Alpha forest cloud sizes in deep space. These observations were to establish whether the correlation with column density is real, further constrain the size of the clouds at two different redshift ranges, and probe cloud sizes on scales up to 1 Mpc. As described in yesterday’s report and HSTAR 7247, the guide star acquisition and re-acquisitions for the first iteration of this proposal failed and the take data flag remained down, affecting four observations. Then, as described in 2.1 and HSTAR 7248, the acquisition and re-acquisitions for the second iteration also failed. The take data flag remained down, affecting seven observations. 1.5 Completed WF/PC-2 8441 (WF/PC-2 Decontaminations and Associated Observations Pt. 1/2) The WF/PC-2 was to be used to make its monthly decontamination operations. As described in yesterday’s report and HSTAR 7247, the guide star acquisition and re-acquisitions for this proposal failed and the take data flag remained down. 1.6 Completed Two Sets of WF/PC-2 8457 (UV Earthflats) The WF/PC-2 was used to make a ultraviolet Earth flat calibration image. This is to improve the quality of pipeline the ultraviolet flatfields. As described in yesterday’s report and HSTAR 7247, the guide star acquisition and re-acquisitions for the first iteration of this proposal failed and the take data flag remained down. Otherwise, the observations were completed as scheduled, and no further problems were reported. 1.7 Completed Two Sets of STIS/CCD 8202 (Bright Quasar Close Lensing Search) The Space Telescope Imaging Spectrograph (CCD) was used to make observations of HCLQ0047-0 and HCLQ2309-5 as part of a second generation survey to search for bright lensing quasars. With the restored PSF, new images will be considerably more sensitive to the presence of both close separation and faint lens components, and because STIS reaches two magnitudes fainter the PC, the new snapshot images will have much greater dynamic range. The observations were executed as scheduled, and no anomalies were noted. 1.8 Completed Two Sets of WF/PC-2 6631 (Imaging a Cluster of Galaxies at Redshift 2.38) The WF/PC-2 was used to make observations of the galactic cluster Gal-Clus-2073 which is located at a distance of z = 2.38. This is a sub-set of about four galaxies within the cluster that forms a very compact {< 100 kpc} galaxy group and may be merging to form a giant elliptical galaxy. The observations were executed as scheduled, and no anomalies were noted. 1.9 Completed STIS/CCD 8406 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal Continuation V) The Space Telescope Imaging Spectrograph (CCD) was used to make low and high galactic latitude observations taken in parallel with the prime science observations as part of a POMS proposal. This is a STIS non-scripted parallel Condition V proposal. The observations were completed as scheduled, and no anomalies were reported. 1.10 Completed WF/PC-2 8274 (A Size Survey of Cometary Nuclei) The WF/PC-2 was used to make observations of the comet Ultra as part of a size survey of cometary nuclei. This is part of a program to determine the basic physical properties of a large fraction of the population of short-period comets. The observations were completed as planned, and no anomalies were reported. 1.11 Completed FGS 8335 (Calibrating Stellar Models with the Pleiades: Resolving the Distance Discrepancy) The Fine Guidance Sensors were used to make astrometric observations of the star H173 (and several reference stars) that are located in the Pleiades cluster. This is the first phase of a multi-year effort to measure true masses for low-mass dwarfs in the Pleiades and to determine an accurate distance to this fundamental cluster. The observations were completed as scheduled, and no anomalies were noted. 1.12 Completed FGS 8464 (Monitor the Long-Term Stability of FGS1R in Position Mode) The Fine Guidance Sensors were used to make astrometric observations of the star OF56 (and several other reference stars) in order to monitor the long-term stability of FGS-1R in the Position Mode. The goal of these measurements is to monitor the internal positional stability of the FGSs by measuring many stars in the M-35 check field repeatedly over a one year period. The observations were completed as planned and no anomalies were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions: 8 Successful: 7 Scheduled Re-acquisitions: 7 Successful: 4 As documented in HSTAR 7248, the acquisition at 251/194124Z failed after acquiring temporarily several times. A large error null bias of 726.672 was present. FHST maps were performed at 251/1956Z and 251/2028Z, followed by attitude reference updates; however, the re-acquisitions at 251/212851Z, 251/225333Z, and 252/003214Z also failed when the scan step limit was exceeded. The proposal described in 1.4 was affected. 2.2 FHST Updates: Scheduled: 10 Successful: 10 2.3 Operations Notes: The SM3A Joint Integrated Simulation (JIS) #5 (EVA1/Replanning/EVA2) is currently being executed. The operations team remains in the SAMS back-up facility. Using ROP SR-1, the SSR EDAC error counter was cleared twice. Due to the acquisition failure mentioned in 2.1, ROP NS-12 was utilized to reset the STIS flight software error counter at 251/2018Z. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH